# Unsourced Random Access: A Recent Paradigm for Massive Connectivity

### v1

@@ V-A Single-Antenna Fading Channels
We first take a look at the existing URA fading solutions assuming a single-antenna receiver. The works in [55, 67, 119, 120, 83, 114, 115, 56, 72, 74, 116, 75, 73] introduce practical coding schemes tailored for the URA scheme on fading channels. Within this collection of works, some assume that the received signals from all users are synchronized [67, 55, 83, 114, 115, 56], while others take into account asynchronous transmission [116, 72, 74, 75, 73]. To estimate the delay of asynchronous signals, [116, 72, 74, 73] use OFDM transmission to convert the time delay to the phase shift in the frequency domain, while the authors in [75] consider time-domain transmission where the delay of a signal is detected via correlation-based approach.
Focusing on the coding schemes in the context of Rayleigh fading channels, we have the following classification:
TIN: Similar to GMAC, the TIN solution for fading channels operates under the assumption of the presence of just a single user and treats other interfering users as noise. Consequently, this approach results in an exceptionally low average signal-to-noise ratio (SNR) due to the fact that it encounters interference-plus-noise as the system’s overall noise level, leading to a very low performance.
@@
Figure 11: Encoding (top) and decoding (bottom) structures of a CCS scheme with 4 slots [55].
CCS [56, 75, 120]: As explained in the previous section, CCS was originally proposed for URA over GMAC. Nevertheless, it has been applied to the fading scenario as well in several works. The encoding and decoding structures of a CCS-based scheme with four slots are shown in Fig. 11, where each user’s signal is divided into four parts, with each one being appended to a parity-check bit. The result is mapped to a codebook to construct the pilot of the user in a slot. Also, at the receiver, sub-messages of different users are detected at each slot, and corresponding ones in different slots are stitched together to form the complete message sequence of a user.
@@
In [56], the authors modify the tree decoder of the CCS scheme in [55] by replacing it with two codes capable of correcting $t$ errors: the first one modifies the conventional tree code by giving it the ability to recover $t$ errors (called $t$-tree), and in the second one, Reed-Solomon (RS) code is employed. In [120], the outer encoded message bits are partitioned into sub-blocks, and each sub-block is transmitted in a different slot after CS encoding, with AMP at the receiver. Moreover, in [75], the authors utilize CCS for the asynchronous URA model. More specifically, during the encoding stage, they append $T_{z}$ additional zeros to the end of each CS sequence, where $T_{z}$ represents the maximum possible delay of the system. At the receiving end, the decoder takes into account all potential delays, ranging from one symbol to a maximum of $T_{z}$-symbol delays when performing CS decoding.
Polar coding with SIC [83]: We next delve into the encoding and decoding procedures of the schemes employing polar codes in conjunction with TIN-SIC. The frame in this scheme is divided into $S$ slots. Each user randomly selects a slot to transmit its polar coded and modulated signal. For decoding, the scheme considers a TIN strategy followed by an SIC. This is because, different users in a slot are prone to different fading coefficients, some of them with larger coefficients are more likely to be decoded by TIN. After decoding the strongest user’s signal while treating others as noise, its channel coefficient is estimated using the orthogonal matching pursuit (OMP) technique. Given the estimated channel coefficient and the decoded signal of a user, its contribution can be removed from the received signal using SIC. This strategy continues until no signal is successfully detected during an iteration. It is important to emphasize that the determination of whether a decoded message is successful relies on the use of a cyclic redundancy check (CRC) message sequence. The encoding and decoding process of the scheme with polar coding with SIC in a slot with $K$ users is demonstrated in Fig. 12.
Figure 12: Encoding and decoding process of the polar code with SIC scheme in [83] in a slot with $K$ users.
There are also some other solutions focusing on URA over fading channels with a single-antenna receiver. For instance, in [119], the active users utilize the entire transmission frame to transmit their data. Specifically, the channel frame is divided into pilot and data parts. The active users pick a pilot sequence from a common codebook for transmission in the pilot part, and they repeat their codeword bits along the rest of the transmission frame with permutation and scrambling in the data part. At the receiver, AMP is employed for joint pilot detection and channel estimation, and the data part is recovered by iterative data estimation and interference cancellation using the repeated codeword bits. Furthermore, the authors of [116] study URA over frequency-selective channels and propose OFDM to mitigate the channel effects combined with CS-based activity detection and TIN-SIC for message recovery at the receiver side.
LDPC with joint decoding [114, 115]: For encoding, these algorithms partition the time frame into distinct slots. Each user transmits its LDPC codeword through a randomly chosen slot. As previously mentioned, slot allocation offers two significant benefits: it reduces the computational load on the decoder by reducing the length of the LDPC signal, and mitigates per-slot interference by distributing users across slots.
@@
The schemes utilize an iterative belief propagation (BP) decoder, as illustrated in Fig. 13. The graph incorporates four distinct node types: variable nodes (in red), and check nodes (in blue), which constitute the traditional LDPC decoder; functional nodes (in green) corresponding to the symbols of the received signal; and a fourth type (in magenta) corresponding to the fading coefficients. This graph allows them to simultaneously update the estimated values of both the channel coefficients and the symbols for each user.
@@
LDPC coding with SIC [72, 73]: This coding design builds upon the schemes presented in [114, 115], with the key distinction of incorporating OFDM techniques to mitigate the effects of the asynchronous transmission by converting the time delay into a phase shift, and employing SIC instead of joint decoding.
@@
The performance of different coding schemes in the single-antenna fading URA is compared in Fig. 14 assuming $n=30000$, $B=100$, and $\epsilon=0.1$. The CCS-based scheme in Andreev et al., 2022 exhibits lower energy efficiency compared to the other ones since it employs a sub-optimal coding scheme, while others use strong channel codes such as LDPC or polar codes. The performance of the LDPC-based scheme with joint decoding in Kowshik et al., 2020 is inferior to that of the ones in Pradhan et al., 2022, Kowshik et al., 2019, Andreev et al., 2020, which can be attributed to the absence of an efficient SIC block in this LDPC-based scheme. On the other hand, the TIN-SIC approach in Kowshik et al., 2019, Andreev et al., 2020 offers a better performance as it allows the receiver to recover the messages of users with a stronger fading coefficient, subtract their effects, and decode the remaining users in subsequent iterations, especially when it is combined with polar coding as in Andreev et al., 2020. Nevertheless, the TIN-SIC scheme in Andreev et al., 2020 is outperformed by the spreading based scheme in Nassaji et al., 2022, and the ODMA-based scheme employing polar codes in Ozates et al., 2024, offering the best performance among the proposed schemes for single-antenna fading case.
A summary of URA studies over fading channels involving a single receive antenna is given in Table III.
Table III: Summary of available coding schemes in URA over fading channels.
Paper Contribution Description Synchron. vs. Asynch.
Kowshik et al., 2019&2021 [84, 113] Bounds Achievable and converse bounds for known and unknown CSI in an asymptotic setting + comparing with CDMA, TDMA, and FDMA. Synchron.
Andreev et al., 2020 [83] Bounds & coding scheme: polar code with SIC Non-asymptotic achievability and converse bounds for unknown CSI + a new coding scheme containing slotting, polar code, and SIC. Synchron.
Kowshik et al., 2019&2020 [114, 115] Bounds& LDPC code with joint decoding Achievable and converse bounds for known and unknown CSI in asymptotic and non-asymptotic manners + a new coding scheme containing LDPC code, slotting, and iterative joint decoding algorithm. Synchron.
Andreev et al., 2022 [56] Coding scheme: CCS-based Improving CCS in [55] by replacing the conventional outer tree code with 1) a code capable of correcting $t$ errors, and 2) a Reed–Solomon code. Synchron.
Yun et al., 2024 [120] Coding scheme: CCS-based A slotted scheme including CCS and AMP in a block fading model. Synchron.
Nassaji et al., 2022 [119] Coding scheme: Spreading-based A non-slotted scheme including, preamble transmission, channel coding, repetition, permutation, scrambling, and SIC. Synchron.
Andreev et al., 2019 [72, 73] Coding scheme: LDPC code with SIC A new coding scheme containing LDPC code, OFDM, slotting, and SIC. Asynch.
Chen et al., 2017 [74] Coding scheme: OFDM with SIC A new coding scheme containing OFDM and SIC. Asynch.
Ozates et al., 2023 [116] Coding scheme: OFDM with SIC A new coding scheme in the frequency-selective channel, containing OFDM, pilot transmission, slotting, polar code, and SIC. Asynch.
Amalladinne et al., 2019 [75] Coding scheme: CCS-based They extend CCS in [55] to accommodate asynchronous scenarios. Asynch.
V-B MIMO Fading Channels
@@
We now turn our focus to the URA solutions for MIMO fading channels. In [87, 121], achievability and converse bounds are obtained for URA over MIMO Rayleigh fading channels, whose ML-based achievability bound is obtained by employing a projection decoder in [122]. Furthermore, several studies have investigated practical coding schemes for the MIMO setting, which can be categorized into four groups: CCS-based, tensor-based, pilot-based, and Bayesian approaches [60, 62, 58, 59, 123, 124, 76, 37, 77, 78, 39, 79, 125, 80, 63, 82, 38], as depicted in Fig. 15.
Figure 15: Coding strategies in URA over MIMO fading channels.
@@
As the computational complexity of mapping bit sequences of each user to a codebook exponentially increases with the message length $B$, CCS-based algorithms are employed in which the bit sequence of each user is divided into $S$ sub-messages [58, 59]. Subsequently, these sub-messages are mapped to pilot codebooks with lower dimensions (see the top part of Fig. 11) to obtain $S$ different pilots. Then, pilots are transmitted through $S$ different slots. At the receiving end, $K_{a}$ sub-messages are detected within each slot, and they are then combined (see the lower part of Fig. 11) to form the complete length-$B$ bit sequence for each user. To establish connections between these sub-messages, the authors in [58] incorporate parity-check bits into each sub-message, while [59] relies on the correlation between channel coefficients of each user in different slots (assuming a quasi-fading channel model). Eliminating the necessity for additional parity bits enhances the performance of [59] in comparison to [58].
@@
Eliminating the parity bits in [58] turns the problem into an uncoupled compressed sensing one, which is also considered in some recent works [123, 124, 126, 127]. To stitch the decoded sub-messages, the authors of [123] employ a joint Bayesian decoding exploiting the channel characteristics, while geometric similarities between the codeword covariance matrices are utilized in the decoding algorithm in [126] and the similarity of the estimated channel vectors are used in [127]. The structure of the inner SPARC code allows parallel single-user outer decoding of device messages in [124], which mitigates the need for the parity check bits. The uncoupled structure is also employed in [128] to design a URA scheme for near-field communications. Furthermore, a two-phase CS coding scheme is proposed in [129], where a long sub-block that is recovered in the first phase is used for channel estimation, and very short lengths of sub-blocks are recovered in the second phase by employing a mutually orthogonal codebook and a hybrid AMP in an uncoupled structure. Moreover, the heterogeneous mobility of the massive number of users is considered in [130] via a two-codebook aided URA scheme with covariance-based maximum likelihood detection (CB-MLD) at the receiver, where the static users employ a larger codebook while the mobile users utilize a smaller codebook in both the codeword length and the number of codewords.
Figure 16: Encoding and decoding structures of tensor-based schemes, where each user divides its message sequence into $d$ parts (here $d=2$), then applies Kronecker product on them [62].
@@
Another trend in MIMO URA is the tensor-based schemes, where each user sends a rank-1 tensor of order $d$. Therefore, the received signal is a rank-$K_{a}$ tensor of order $d+1$ (considering the channel coefficient vector as an extra dimension of the tensor) [60, 61, 62, 63]. For decoding, they conduct tensor decomposition on the received signal, followed by de-quantization of the tensors. In particular, in [60], by tensor decomposition, the rank-$K_{a}$ tensor of order $d+1$ is decomposed into $K_{a}$ rank-1 tensors of order $d+1$, which is equivalent to separating the signals of $K_{a}$ users, where the transmitted tensors can be identified by the first $d$ dimensions, and the $(d+1)$th dimension gives the channel vector of the user. Then, by de-quantizing the tensor of each user, its transmitted bit-sequence is detected. This scheme is extended to the case with the presence of timing offsets in [61] by employing OFDM. On the other hand, the authors of [62] propose a block-term decomposition for the decomposition of the 3-dimensional tensors, where two low-rank factor matrices and a factor vector are utilized in each block term. The overall coding structure of tensor-based methods is depicted in Fig. 16.

### v2

@@ V-B Single-Antenna Fading Channels
We first take a look at the existing URA fading solutions that assume a single-antenna receiver. The works in [72, 94, 96, 85, 80, 89, 90, 92, 95, 93, 91] introduce practical coding schemes tailored for the URA scheme on single-antenna fading channels. Within this collection of works, some assume that the received signals from all users are synchronized [72, 85, 80, 89], while others take into account asynchronous transmission [95, 90, 92, 93, 91]. To estimate the delay of asynchronous signals, [95, 90, 92, 91] use OFDM transmission to convert the time delay to the phase shift in the frequency domain, while the authors in [93] consider time-domain transmission where the delay of a signal is detected via correlation-based approach.
Focusing on the Rayleigh fading channels, in the following, we present the existing fading URA solutions, categorized based on their general approach.
@@
Figure 10: Another representation of encoding (top) and decoding (bottom) structures of a CCS scheme with 4 slots [60].
V-B1 CCS-based schemes
@@
As explained in the previous section, CCS was originally proposed for URA over GMAC. Nevertheless, it has been applied to the fading scenario as well in several works [89, 93, 96]. In Fig. 10, another representation of the encoding (upper half of the figure) and decoding (lower half) structures of a CCS-based scheme with four slots are depicted, where each user’s signal is divided into four parts, with each one being appended to a parity-check bit, which is then mapped to a codebook to construct the pilot of the user in a slot. Also, at the receiver, sub-messages of different users are detected at each slot, and corresponding ones in different slots are stitched together to form the complete message sequence of a user. In [89], the authors modify the tree decoder of the GMAC CCS scheme in [60] by replacing it with two codes capable of correcting $t$ errors, in which the first one modifies the conventional tree code by giving it the ability to recover $t$ errors (called $t$-tree), and the second one is a Reed-Solomon (RS) code. In [96], the outer encoded message bits are partitioned into sub-blocks, and each sub-block is transmitted in a different slot after CS encoding, with AMP at the receiver. Moreover, in [93], the authors utilize CCS for the asynchronous URA model. More specifically, during the encoding stage, they append $T_{z}$ additional zeros to the end of each CS sequence, where $T_{z}$ represents the maximum possible delay of the system. At the receiving end, the decoder takes into account all potential delays, ranging from one symbol to a maximum of $T_{z}$-symbol delays when performing the CS decoding.
V-B2 Joint decoding-based schemes
@@
In [80], a joint decoding approach is employed. For encoding, this scheme partitions the time frame into distinct slots. Each user transmits its LDPC codeword through a randomly chosen slot. As previously mentioned, slot allocation offers two significant benefits: it reduces the computational load on the decoder by reducing the length of the LDPC signal, and mitigates per-slot interference by distributing users across slots. This scheme utilizes an iterative belief propagation (BP) decoder, as illustrated in Fig. 11. The graph incorporates four distinct node types: variable nodes (in red), and check nodes (in blue), which constitute the traditional LDPC decoder; functional nodes (in green) corresponding to the symbols of the received signal; and a fourth type (in magenta) corresponding to the fading coefficients. This graph allows them to simultaneously update the estimated values of both the channel coefficients and the symbols for each user.
@@
V-B3 TIN-SIC-based schemes
A lower-complexity alternative to joint decoding is TIN-based decoding. Similar to GMAC, the TIN solution for fading channels operates under the assumption of the presence of just a single user and treats other interfering users as noise. Consequently, this approach results in an exceptionally low average SNR due to the fact that it encounters interference-plus-noise as the system’s overall noise level, leading to a very low performance. However, when combined with an appropriate channel code and SIC, it can lead to good performance.
Table IV: A summary of URA solutions for fading channels with a single-antenna receiver.
Paper Approach Description Synchron. vs. Asynch.
Kowshik et al., 2019&2021 [83, 84] - Achievable and converse bounds for known and unknown CSI in an asymptotic setting + comparing with CDMA, TDMA, and FDMA. Synchron.
Andreev et al., 2019 [90, 91] TIN-SIC A new coding scheme containing LDPC code, OFDM, slotting, and SIC. Asynch.
Andreev et al., 2020 [85] TIN-SIC Non-asymptotic achievability and converse bounds for unknown CSI + a new coding scheme containing slotting, polar code, and SIC. Synchron.
Kowshik et al., 2019&2020 [80] joint decoding Achievability and converse bounds for known and unknown CSI + a scheme containing LDPC code, slotting, and iterative joint decoding. Synchron.
Amalladinne et al., 2019 [93] CCS-based They extend CCS in [60] to accommodate asynchronous scenarios. Asynch.
Andreev et al., 2022 [89] CCS-based Improving CCS in [60] by replacing the conventional outer tree code with 1) a code capable of correcting $t$ errors, and 2) a Reed–Solomon code. Synchron.
Yun et al., 2024 [96] CCS-based A slotted scheme including CCS and AMP in a block fading model. Synchron.
Nassaji et al., 2022 [94] Spreading-based A non-slotted scheme including, preamble transmission, channel coding, repetition, permutation, scrambling, and SIC. Synchron.
Chen et al., 2017 [92] OFDM+ SIC A new coding scheme containing OFDM and SIC. Asynch.
Ozates et al., 2023 [95] OFDM+SIC A new coding scheme in the frequency-selective channel, containing OFDM, pilot transmission, slotting, polar code, and SIC. Asynch.
TIN-SIC with LDPC coding: In [90], building upon their asynchronous schemes presented in [80], the authors present a low-complexity coding scheme which incorporates orthogonal frequency division multiplexing (OFDM) techniques to mitigate the effects of the asynchronous transmission by converting the time delays into phase shifts, and employing SIC to decode the user packets. More numerical results for this scheme is presented in [91].
TIN-SIC with Polar Coding: In [85], a TIN-SIC solution in conjunction with polar coding is presented, where the time frame in this approach is divided into $S$ slots. Each user randomly selects a slot to transmit its polar coded and modulated signal. For decoding, the scheme considers a TIN strategy followed by an SIC. This is because different users in a slot are prone to different fading coefficients; some of them with larger coefficients are more likely to be decoded by TIN. After decoding the strongest user’s signal while treating others as noise, its channel coefficient is estimated using the orthogonal matching pursuit (OMP) technique. Given the estimated channel coefficient and the decoded signal of a user, its contribution can be removed from the received signal using SIC. This strategy continues until no signal is successfully detected during an iteration. It is important to emphasize that the determination of whether a decoded message is successful relies on the use of a cyclic redundancy check (CRC) message sequence. The encoding and decoding process of the scheme with polar coding with SIC in a slot with $K$ users is demonstrated in Fig. 12.
Figure 12: Encoding and decoding processes of TIN-SIC scheme with polar codes in [85] in a slot with $K$ users.
V-B4 Alternative solutions
There are also some other solutions focusing on URA over fading channels with a single-antenna receiver. For instance, in [94], the active users utilize the entire transmission time frame to transmit their data. Specifically, the time frame is divided into pilot and data parts. The active users pick a pilot sequence from a common codebook for transmission in the pilot part, and they repeat their codeword bits along the rest of the transmission time frame with permutation and scrambling in the data part. At the receiver, AMP is employed for joint pilot detection and channel estimation, and the data part is recovered by iterative data estimation and interference cancellation using the repeated codeword bits. Furthermore, the authors of [95] study URA over frequency-selective channels and propose OFDM to mitigate adverse channel effects combined with CS-based activity detection and TIN-SIC for message recovery at the receiver side.
@@ V-B4 Alternative solutions
Summary and Performance Comparison:
@@
A summary of URA studies over fading channels involving a single-antenna receiver is given in Table IV. Moreover, the performance of different coding schemes for the single-antenna fading URA setup is compared in Fig. 13 assuming $n=30000$, $B=100$, and $\epsilon=0.1$. The reason that some diagrams seem not to be fully plotted is that they diverge after a certain active user load, i.e., they cannot achieve the target PUPE no matter how much the $E_{b}/N_{0}$ is increased. As can be observed from Fig. 13, the CCS-based scheme in [89] exhibits lower energy efficiency compared to the other ones since it employs a sub-optimal coding scheme, while others use strong channel codes such as LDPC or polar codes. The performance of the LDPC-based scheme with joint decoding in [80] is inferior to that of the ones in [72], [91] and [85], which can be attributed to the absence of an efficient SIC block in this LDPC-based scheme. On the other hand, the TIN-SIC approach in [91] and [85] offers a better performance as it allows the receiver to recover the messages of users with a stronger fading coefficient, subtract their effects, and decode the remaining users in subsequent iterations, especially when it is combined with polar coding as in [85]. Nevertheless, the TIN-SIC scheme in [85] is outperformed by the spreading-based scheme in [94], and the ODMA-based scheme employing polar codes in [78], which offers the best performance among the proposed schemes, due to the sparsity of the transmissions.
Computational Complexity:
We now compare the computational complexity of the single-antenna fading URA solutions described in this subsection. The tree-code-based CCS approach has a low complexity, which is linear in the list of messages at each decoding step and exponential in the number of message bits in each slot [89]. In addition, the decoding operations in [89] can be implemented by additions, which are easier to implement in hardware compared to multiplications. The complexity of the $T$-fold ALOHA-based schemes employing LDPC with joint decoding [80] or polar coding with TIN-SIC [85] is mainly determined by $T$, that is, the number of users that the decoder tries to recover in each slot. Note that TIN-SIC is more preferable than joint decoding since it utilizes single-user decoding, which makes it generally less complex and more parallelizable. Spreading-based and ODMA-based schemes, proposed in [94] and [78], respectively, have higher complexities compared to their counterparts in exchange for their superior performance. For instance, the complexity of MMSE filtering for fading coefficient estimation in [78] is cubic in $K_{a}$.
V-C MIMO Fading Channels
@@
We now turn our focus to the URA solutions for MIMO fading channels. Several studies have investigated practical coding schemes for the MIMO URA setup, which can be categorized into five groups: CCS-based, tensor-based, pilot-based, ODMA-based, and Bayesian approaches, with inevitable overlaps between them. The MIMO URA solutions based on these approaches are reviewed in the rest of this subsection.
Figure 14: Encoding and decoding structures of tensor-based schemes for $d=2$ [109].
V-C1 CCS-based schemes
@@
In CCS-based algorithms for MIMO fading URA [97, 98], the bit sequence of each user is divided into $S$ sub-messages, which are then mapped to codewords in pilot codebooks with lower dimensions to obtain $S$ different pilots, transmitted through $S$ different slots. At the receiving end, $K_{a}$ sub-messages are detected within each slot, and they are then combined to form the complete length-$B$ bit sequence for each user. To establish connections between these sub-messages, the authors in [97] incorporate parity-check bits into each sub-message, while [98] relies on the correlation between channel coefficients of each user in different slots (assuming a quasi-fading channel model). Eliminating the necessity for additional parity bits enhances the performance of [98] in comparison to [97].
@@
Eliminating the parity bits in [97] turns the problem into an uncoupled compressed sensing one, which is also considered in some recent works [99, 102, 100, 101]. To stitch the decoded sub-messages, the authors of [99] employ a joint Bayesian decoding exploiting the channel characteristics, while geometric similarities between the codeword covariance matrices are utilized in the decoding algorithm in [100], and the similarity of the estimated channel vectors are used in [101]. The structure of the inner SPARC code allows parallel single-user outer decoding of user messages in [102], which mitigates the need for the parity check bits. An uncoupled structure is employed in [103] to design a CCS-based URA scheme for near-field communications. Furthermore, a two-phase CS coding scheme is proposed in [104], where a long sub-block recovered in the first phase is used for channel estimation, and very short lengths of sub-blocks are recovered in the second phase by employing a mutually orthogonal codebook and a hybrid AMP in an uncoupled structure. Moreover, the heterogeneous mobility of the massive number of users is considered in [105] via a two-codebook aided URA scheme with covariance-based maximum likelihood detection (CB-MLD) at the receiver, where the static users employ a larger codebook while the mobile users utilize a smaller one in both the codeword length and the number of codewords. Moreover, an uncoupled CS scheme employing polar coding and power diversity in combination with AMP-based decoding and SIC at the receiver side is proposed in [106].
V-C2 Tensor-based schemes
@@
Another trend in MIMO URA is the tensor-based schemes [107, 108, 109], where each user sends a rank-1 tensor of order $d$. Therefore, the received signal is a rank-$K_{a}$ tensor of order $d+1$ (considering the channel coefficient vector as an extra dimension of the tensor). For decoding, they conduct tensor decomposition on the received signal, followed by de-quantization of the tensors. In particular, in [107], by tensor decomposition, the rank-$K_{a}$ tensor of order $d+1$ is decomposed into $K_{a}$ rank-1 tensors of order $d+1$, which is equivalent to separating the signals of $K_{a}$ users, where the transmitted tensors can be identified by the first $d$ dimensions, and the $(d+1)$th dimension gives the channel vector of the user. Then, by de-quantizing the tensor of each user, its transmitted bit-sequence is detected. By employing OFDM transmission, this scheme is extended to the case with timing offsets in [108]. On the other hand, the authors of [109] propose a block-term decomposition for the decomposition of 3-dimensional tensors, where two low-rank factor matrices and a factor vector are utilized in each block term. The overall coding structure of tensor-based methods is depicted in Fig. 14.
@@
V-C3 Pilot-based schemes
Pilot-based techniques constitute another category of algorithms in the context of MIMO URA systems, where, the idea is to divide user bit sequences into pilot and data parts. In the existing solutions, the transmitted signal of each user also consists of two parts, as depicted in Fig. 15: the pilot part which is obtained by mapping the pilot bits to a codeword in a pilot codebook (a non-orthogonal codebook [111, 12, 113, 112, 14, 110], or a codebook consisting of multiple stages of orthogonal pilots [13]), and the data part which is encoded by a channel code (a polar [13, 111, 12, 14] or an LDPC code [113, 112]). For decoding, as shown in Fig. 17, these schemes iteratively employ the following steps:
  • At the initial stage, the receiver detects the active pilots using a user activity detection technique, such as generalized OMP (gOMP) [14, 110], AMP [112, 114], multiple-measurement vector AMP (MMV-AMP) [111], energy detector [12], NP hypothesis testing [13] or CB-MLD [113].
  • Using the estimated channel coefficients, soft estimates of the transmitted codewords are obtained via decoupling techniques, such as MRC [111, 13, 14, 110], MMSE [12] or AMP [113], which are then used for generating log-likelihood ratio (LLR) values to be fed to single-user decoders [13, 111, 12, 113, 14, 114] or a joint factor graph [112].
  • In most pilot-based schemes, following the SIC approach, the contribution of the successfully decoded codewords is removed from the received signal to decrease the interference level at the next iterations. Note that in [13, 12, 14, 110], the channel vectors are re-estimated using the successfully decoded codewords to have a better interference cancellation at the SIC stage.